OpenStudy (anonymous):

Evaluate the definite integrals [tex]\int_{{0}_{1}}{\frac{dr}{((7-5r)^{2})^{1/3}}}[/tex]

OpenStudy (anonymous):

looks horrid this way :)

OpenStudy (anonymous):

ok this looks like gibberish one sec

myininaya (myininaya):

\[\int\limits_{0}^{1}\frac{dr}{(7-5r)^\frac{2}{3}}\]

OpenStudy (anonymous):

This is easier than the last one

myininaya (myininaya):

let u=7-5r du=-5 dr -1/5 du=dr if x=1, then u=7-5(1)=2 if x=0, then u=7-5(0)=7

OpenStudy (anonymous):

myininaya dont write the solution, please. Purplec should be able to solve this, we can help if he is stuck. I was too slow :DDD

myininaya (myininaya):

ok andras

OpenStudy (anonymous):

do you know where you got stuck with this?

OpenStudy (anonymous):

well the solution is there now, I hope it is clear. For many integrals substitution is a good solution. The only thing you have to be carefull how the dx changes to du

OpenStudy (anonymous):

omg...would you believe all that time i was fighting to put in in the post and it delted!!! :@

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

myinaya put in right but i forgot about my algerba that i cant raise it 2/3

OpenStudy (anonymous):

and yes it is easier lol

OpenStudy (anonymous):

integration by parts is the other tool that you have to use a lot

OpenStudy (anonymous):

for example for integral x^2 * sinx

OpenStudy (anonymous):

i am not a he i am a she thank you very much...why would a guycall himself purple lol? newho

OpenStudy (anonymous):

okok :) I am a guy and I do like purple

myininaya (myininaya):

also partial fractions is good tool to integrate